i dont understand how we find the domain without graphing our problem is Find the range of the function for the domain {–3, 1, 4.5}.
y = –x + 1 but i dont understand

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

In its simplest form, the domain is all the values that go into a function, and the range is all the values that come out.
So if you input {–3, 1, 4.5} for x what will you get out for y

Answer by stanbon(75887)

i dont understand how we find the domain without graphing our problem is Find the range of the function for the domain {–3, 1, 4.5}.
y = –x + 1 but i dont understand
--------
f(-3) = -(-3)+1 = 3 + 1 = 4
f(1) = -1+1 = 0
f(4.5) = -4.5 + 1 = -3.5
---------
Range:: {4,0,-3.5}
-------
Comment:
Finding the domain.
1st: Assume the domain is All Real Numbers.
---
2nd: Look for elements of the Rule which would restrict the domain, such as::
a) even roots of a variable expression; it must be >= zero
b) a variable expression in a denominator; it cannot equal zero
c) log or ln of a variable expression; it must be >= zero
-------------
If the Rule you have does not have a or b or c,
the domain is All Real Numbers.
==============================
Cheers,
Stan H.
==============================